# Chains of model structures arising from modules of finite Gorenstein dimension

### v3

@@ 4.1. Equalities on finitistic dimensions
For any ring $R$, Maaouy [22, Theorem A] proves that $(\mathcal{GF}_{n},\ \mathcal{F}_{n}^{\perp}\cap\mathcal{PGF}^{\perp})$ is a complete and hereditary cotorsion pair, and that $(\mathcal{GF}_{n},\ \mathcal{F}_{n}^{\perp},\ \mathcal{PGF}^{\perp})$ is a hereditary Hovey triple, in $R$-Mod.
It seems to be natural to ask whether $\mathcal{PGF}^{<\infty}$, $\mathcal{GF}^{<\infty}$ and $\mathcal{GF}^{<\infty}$ could be the left parts of cotorsion pairs in $R$-Mod. This problem reasonably leads to several finitistic dimensions.
@@
The (left) finitistic Gorenstein projective dimension ${\rm FGpd}={\rm FGpd}(R)$ ([17, 2.27]) is defined as
@@
The (left) finitistic Gorenstein flat dimension FGfd = FGfd$(R)$ ([17, 3.24]) is defined as
@@
Using an argument of H. Holm, one can prove the following analogue of ([17, Theorem 2.28]).
@@
(1)  ${\rm FPGFd}={\rm FGpd}={\rm Fpd}$.
(2)  If $R$ is right coherent and ${\rm fd}M<\infty$, then ${\rm Gfd}M={\rm fd}M$.
(3)  If $R$ is right coherent, then ${\rm FGfd}={\rm Ffd}\leq{\rm FPGFd}={\rm FGpd}={\rm Fpd}.$
@@
(1)  [17, Theorem 2.28] claims that ${\rm FGpd}={\rm Fpd}$. Thus, it suffices to prove ${\rm FPGFd}={\rm Fpd}$.
@@
If ${\rm pd}M<\infty$, then ${\rm Gpd}M={\rm pd}M$ (see [17, Proposition 2.27]), and hence ${\rm PGFd}M={\rm pd}M$ (see also [12, Corollary 3.7(i)]). It follows that ${\rm Fpd}\leq{\rm FPGFd}.$
@@
(2)   Assume that ${\rm fd}M=m<\infty$. Since $\mathcal{F}\subseteq\mathcal{GF}$, it follows that ${\rm Gfd}M\leq{\rm fd}M=m$. To prove ${\rm Gfd}M=m$, by [17, Theorem 3.14] it suffices to find an injective right $R$-module $I$ such that ${\rm Tor}^{R}_{m}(I,M)\neq 0$. Since ${\rm fd}M=m$, it follows that there is a right $R$-module $N$ such that ${\rm Tor}^{R}_{m}(N,M)\neq 0$. Consider an exact sequence $0\longrightarrow N\longrightarrow I\longrightarrow C\longrightarrow 0$ such that $I$ is an injective right $R$-module. Then by the exact sequence
${\rm Tor}^{R}_{m+1}(C,M)=0\longrightarrow{\rm Tor}^{R}_{m}(N,M)\longrightarrow{\rm Tor}^{R}_{m}(I,M)\longrightarrow{\rm Tor}^{R}_{m}(C,M)$
one sees that ${\rm Tor}^{R}_{m}(I,M)\neq 0$. We are done.
(3)  [17, Theorem 3.24] already claims that if $R$ is right coherent, then ${\rm FGfd}={\rm Ffd}$.
Since $\mathcal{F}\subseteq\mathcal{PGF}\subseteq\mathcal{GF}$, it follows that ${\rm Gfd}M\leq{\rm PGFd}M\leq{\rm fd}M$ for any $R$-module.
If ${\rm fd}M<\infty$, then ${\rm Gfd}M={\rm fd}M$, by (2). It follows that ${\rm PGFd}M={\rm fd}M$. This implies that ${\rm Ffd}\leq{\rm FPGFd}.$ Together with (1) one gets ${\rm FGfd}={\rm Ffd}\leq{\rm FPGFd}={\rm FGpd}={\rm Fpd}.$
@@
(3)  If $R$ is right coherent, then  $(\mathcal{GF}^{<\infty},\ (\mathcal{F}^{<\infty})^{\perp}\cap\mathcal{PGF}^{\perp})=(\mathcal{GF}_{t},\ \mathcal{F}_{t}^{\perp}\cap\mathcal{PGF}^{\perp})$ is a complete and hereditary cotorsion pair, and  $(\mathcal{GF}^{<\infty},\ (\mathcal{F}^{<\infty})^{\perp},\ \mathcal{PGF}^{\perp})=(\mathcal{GF}_{t},\ \mathcal{F}_{t}^{\perp},\ \mathcal{PGF}^{\perp})$ is a hereditary Hovey triple, in $R\mbox{-}{\rm Mod}.$
@@
(1)  By Proposition 4.1, one has ${\rm FGFd}={\rm Ffd}\leq{\rm FPGFd}={\rm FGpd}=t$, thus $\mathcal{PGF}^{<\infty}=\mathcal{PGF}_{t}$ and $\mathcal{P}^{<\infty}=\mathcal{P}_{t}$,  $\mathcal{GP}^{<\infty}=\mathcal{GP}_{t}$  and $\mathcal{P}^{<\infty}=\mathcal{P}_{t}$,  and $\mathcal{GF}^{<\infty}=\mathcal{GF}_{t}$  and $\mathcal{F}^{<\infty}=\mathcal{F}_{t}$. Then the assertions follow from Corollary 2.6, Theorem 3.4, and [22, Theorem A], respectively.∎
@@ 5.1. Induced complete cotorsion pairs in exact categories
$(2)$  Since $\mathcal{X}$ is closed under the kernels of epimorphisms in $\mathcal{A}$, $\mathcal{X}$ is closed under the kernels of deflations in $\mathcal{B}$. Since $(\mathcal{X},\ \mathcal{Y}\cap\mathcal{B})$ is a complete cotorsion pair in $\mathcal{B},$ it follows that it is also hereditary (cf. [29, Lemma 6.17]; or Subsection 1.2). ∎
@@ 5.2. Application 1: Exact model structures on $\mathcal{PGF}_{n}$
Since $K\in\mathcal{P}_{n-1}$ and $Y\in\mathcal{P}_{n}$, it follows that $L\in\mathcal{P}_{n}$. Since $X\in\mathcal{P}_{n}$ and $L\in\mathcal{P}_{n}$, it follows that ${\rm pd}G<\infty.$ While $G\in\mathcal{PGF}$, $G$ is Gorenstein projective, and hence $G$ has to be projective (cf. [11, Proposition 10.2.3]). Thus $Z\in\mathcal{P}_{n}$. ∎
@@ 6.1. Application 3: Exact model structures on $\mathcal{GP}_{n}$
Note that $\mathcal{GP}_{n}$ is closed under extensions, the kernels of epimorphisms, and direct summands (cf. [17, Theorem 2.20, Proposition 2.19]). In particular, $\mathcal{GP}_{n}$ is a weakly idempotent complete exact category.
@@ 6.4. Exact model structures on $\mathcal{GP}^{<\infty}$
It is clear that $\mathcal{P}_{m}\cap\mathcal{GP}^{<\infty}=\mathcal{P}_{m}$. Also note that $\mathcal{GP}_{m}\cap\mathcal{P}^{<\infty}=\mathcal{P}_{m}$ (cf. [17, Proposition 2.27]). Similar to Corollary 6.1 and Theorem 6.3 one has
@@ 7.1. Application 4: Exact model structures on $\mathcal{GF}_{n}$
The following result is due to R. El Maaouy [22]. It is again a theorem of Šaroch - Št́ovíček type; and when $n=0$ it is given in [28, Theorem 4.11].
@@
([22, Theorem 3.4])  Let $R$ be a ring, $M$ an $R$-module, and $n$ a non-negative integer. Then $\mathcal{PGF}^{\perp}\cap\mathcal{GF}_{n}=\mathcal{F}_{n};$ and the following are equivalent::::
@@
The careful reader will find that the statement (2) is slightly stronger than [22, Theorem 3.4]: here $K\in\mathcal{F}_{n-1}$ if $n\geq 1$, rather than $K\in\mathcal{F}_{n}$ for all $n\geq 0$ as in [22, Theorem 3.4]; also, $X\in\mathcal{F}_{n}^{\perp}$ in (2), rather than $X\in\mathcal{F}_{n}\cap\mathcal{F}_{n}^{\perp}$ as in [22, Theorem 3.4]. In this way, Theorem 7.1 coincides with the corresponding versions for modules in $\mathcal{PGF}_{n}$ as in Theorem 2.5, and for modules $\mathcal{GP}_{n}$ as in Theorem 3.3. We stress that the proof of [22, Theorem 3.4] completely holds for Theorem 7.1, thus we will not include a proof.
When $n=0$, the following result is [28, Corollary 4.12].
@@
([22, Theorems A, B])  Let $R$ be a ring and $n$ a non-negative integer. Then
@@
([23, Theorem 3.4(2)])  Let $R$ be a ring, and $n$ a non-negative integer. Then $(\mathcal{F}_{n},\ \mathcal{F}_{n}^{\perp})$ is a complete and hereditary cotorsion pair in $R$-Mod, and cogenerated by a set.
@@
The assertion (2) is implicitly contained in [8]. For convenience we include a justification. By [8, Proposition 2.5(ii)], ${\rm pd}M={\rm fd}M<\infty$. Since $M\in\mathcal{PGF}$, it follows that $M$ is Gorenstein projective ([28, Theorem 4.4]), with finite projective dimension, and hence $M$ is projective (cf. [11, Proposition 10.2.3]). An alternative proof: by Theorem 7.1 one has $M\in\mathcal{PGF}\cap\mathcal{F}_{n}=\mathcal{PGF}\cap\mathcal{GF}_{n}\cap\mathcal{PGF}^{\perp}=(\mathcal{PGF}\cap\mathcal{PGF}^{\perp})\cap\mathcal{GF}_{n}=\mathcal{P}\cap\mathcal{GF}_{n}=\mathcal{P}.$

### v4

@@ 4.1. Equalities on finitistic dimensions
For any ring $R$, Maaouy [25, Theorem A] proves that $(\mathcal{GF}_{n},\ \mathcal{F}_{n}^{\perp}\cap\mathcal{PGF}^{\perp})$ is a complete and hereditary cotorsion pair, and that $(\mathcal{GF}_{n},\ \mathcal{F}_{n}^{\perp},\ \mathcal{PGF}^{\perp})$ is a hereditary Hovey triple, in $R$-Mod.
It seems to be natural to ask whether $\mathcal{PGF}^{<\infty}$, $\mathcal{GP}^{<\infty}$ and $\mathcal{GF}^{<\infty}$ could be the left parts of cotorsion pairs in $R$-Mod. This problem reasonably leads to several finitistic dimensions.
@@
The (left) finitistic Gorenstein projective dimension ${\rm FGpd}={\rm FGpd}(R)$ ([19, 2.27]) is defined as
@@
The (left) finitistic Gorenstein flat dimension FGfd = FGfd$(R)$ ([19, 3.24]) is defined as
@@
(1)  One has ${\rm FPGFd}={\rm FGpd}={\rm Fpd}$.
(2)  If ${\rm fd}M<\infty$, then ${\rm Gfd}M={\rm fd}M$.
(3)  One has ${\rm FGfd}={\rm Ffd}\leq{\rm FPGFd}={\rm FGpd}={\rm Fpd}.$
@@
(1)  [19, Theorem 2.28] claims that ${\rm FGpd}={\rm Fpd}$. Thus, it suffices to prove ${\rm FPGFd}={\rm Fpd}$.
@@
If ${\rm pd}M<\infty$, then ${\rm Gpd}M={\rm pd}M$ (see [19, Proposition 2.27]), and hence ${\rm PGFd}M={\rm pd}M$ (see also [13, Corollary 3.7(i)]). It follows that ${\rm Fpd}\leq{\rm FPGFd}.$
@@
(2)  According to [10, Proposition 3.8], if $M\in\mathcal{PGF}^{\perp}$ then ${\rm Gfd}M={\rm fd}M$. Thus, it suffices to say that ${\mathcal{F}}^{<\infty}\subseteq\mathcal{PGF}^{\perp}$. By [32, Theorem 4.4 or Theorem 4.11], ${\mathcal{F}}\subseteq{\rm PGF}^{\perp}$. Then ${\mathcal{F}}^{<\infty}\subseteq\mathcal{PGF}^{\perp}$ since $\mathcal{PGF}^{\perp}$ is thick.
(3)  I. Emmanouil [10, Proposition 2.3] already claims that ${\rm FGfd}={\rm Ffd}$.
By C. U. Jensen [23, Proposition 6], if ${\rm Fpd}<\infty$ then all flat modules have finite projective dimension. It follows that ${\rm Ffd}\leq\mathcal{\rm Fpd}$.
Together with (1) one gets ${\rm FGfd}={\rm Ffd}\leq{\rm FPGFd}={\rm FGpd}={\rm Fpd}.$
Remark 4.2.
We thank Ioannis Emmanouil for pointing out an error in the original proof of Proposition 4.1(3), and for pointing out to us the paper [23].
@@
(3)  The pair  $(\mathcal{GF}^{<\infty},\ (\mathcal{F}^{<\infty})^{\perp}\cap\mathcal{PGF}^{\perp})=(\mathcal{GF}_{t},\ \mathcal{F}_{t}^{\perp}\cap\mathcal{PGF}^{\perp})$ is a complete and hereditary cotorsion pair, and  $(\mathcal{GF}^{<\infty},\ (\mathcal{F}^{<\infty})^{\perp},\ \mathcal{PGF}^{\perp})=(\mathcal{GF}_{t},\ \mathcal{F}_{t}^{\perp},\ \mathcal{PGF}^{\perp})$ is a hereditary Hovey triple, in $R\mbox{-}{\rm Mod}.$
@@
By Proposition 4.1(3) one has ${\rm FGFd}={\rm Ffd}\leq{\rm FPGFd}={\rm FGpd}=t$. Thus $\mathcal{PGF}^{<\infty}=\mathcal{PGF}_{t}$ and $\mathcal{P}^{<\infty}=\mathcal{P}_{t}$,  $\mathcal{GP}^{<\infty}=\mathcal{GP}_{t}$  and $\mathcal{P}^{<\infty}=\mathcal{P}_{t}$,  and $\mathcal{GF}^{<\infty}=\mathcal{GF}_{t}$  and $\mathcal{F}^{<\infty}=\mathcal{F}_{t}$. Then the assertions (1), (2) and (3) follow from Corollary 2.6, Theorem 3.4, and [25, Theorem A], respectively.∎
@@ 5.1. Induced complete cotorsion pairs in exact categories
$(2)$  Since $\mathcal{X}$ is closed under the kernels of epimorphisms in $\mathcal{A}$, $\mathcal{X}$ is closed under the kernels of deflations in $\mathcal{B}$. Since $(\mathcal{X},\ \mathcal{Y}\cap\mathcal{B})$ is a complete cotorsion pair in $\mathcal{B},$ it follows that it is also hereditary (cf. [33, Lemma 6.17]; or Subsection 1.2). ∎
@@ 5.2. Application 1: Exact model structures on $\mathcal{PGF}_{n}$
Since $K\in\mathcal{P}_{n-1}$ and $Y\in\mathcal{P}_{n}$, it follows that $L\in\mathcal{P}_{n}$. Since $X\in\mathcal{P}_{n}$ and $L\in\mathcal{P}_{n}$, it follows that ${\rm pd}G<\infty.$ While $G\in\mathcal{PGF}$, $G$ is Gorenstein projective, and hence $G$ has to be projective (cf. [12, Proposition 10.2.3]). Thus $Z\in\mathcal{P}_{n}$. ∎
@@ 6.1. Application 3: Exact model structures on $\mathcal{GP}_{n}$
Note that $\mathcal{GP}_{n}$ is closed under extensions, the kernels of epimorphisms, and direct summands (cf. [19, Theorem 2.20, Proposition 2.19]). In particular, $\mathcal{GP}_{n}$ is a weakly idempotent complete exact category.
@@ 6.4. Exact model structures on $\mathcal{GP}^{<\infty}$
It is clear that $\mathcal{P}_{m}\cap\mathcal{GP}^{<\infty}=\mathcal{P}_{m}$. Also note that $\mathcal{GP}_{m}\cap\mathcal{P}^{<\infty}=\mathcal{P}_{m}$ (cf. [19, Proposition 2.27]). Similar to Corollary 6.1 and Theorem 6.3 one has
@@ 7.1. Application 4: Exact model structures on $\mathcal{GF}_{n}$
The following result is due to R. El Maaouy [25]. It is again a theorem of Šaroch - Št́ovíček type; and when $n=0$ it is given in [32, Theorem 4.11].
@@
([25, Theorem 3.4])  Let $R$ be a ring, $M$ an $R$-module, and $n$ a non-negative integer. Then $\mathcal{PGF}^{\perp}\cap\mathcal{GF}_{n}=\mathcal{F}_{n};$ and the following are equivalent::::
@@
The careful reader will find that the statement (2) is slightly stronger than [25, Theorem 3.4]: here $K\in\mathcal{F}_{n-1}$ if $n\geq 1$, rather than $K\in\mathcal{F}_{n}$ for all $n\geq 0$ as in [25, Theorem 3.4]; also, $X\in\mathcal{F}_{n}^{\perp}$ in (2), rather than $X\in\mathcal{F}_{n}\cap\mathcal{F}_{n}^{\perp}$ as in [25, Theorem 3.4]. In this way, Theorem 7.1 coincides with the corresponding versions for modules in $\mathcal{PGF}_{n}$ as in Theorem 2.5, and for modules $\mathcal{GP}_{n}$ as in Theorem 3.3. We stress that the proof of [25, Theorem 3.4] completely holds for Theorem 7.1, thus we will not include a proof.
When $n=0$, the following result is [32, Corollary 4.12].
@@
([25, Theorems A, B])  Let $R$ be a ring and $n$ a non-negative integer. Then
@@
([26, Theorem 3.4(2)])  Let $R$ be a ring, and $n$ a non-negative integer. Then $(\mathcal{F}_{n},\ \mathcal{F}_{n}^{\perp})$ is a complete and hereditary cotorsion pair in $R$-Mod, and cogenerated by a set.
@@
The assertion (2) is implicitly contained in [8]. For convenience we include a justification. By [8, Proposition 2.5(ii)], ${\rm pd}M={\rm fd}M<\infty$. Since $M\in\mathcal{PGF}$, it follows that $M$ is Gorenstein projective ([32, Theorem 4.4]), with finite projective dimension, and hence $M$ is projective (cf. [12, Proposition 10.2.3]). An alternative proof: by Theorem 7.1 one has $M\in\mathcal{PGF}\cap\mathcal{F}_{n}=\mathcal{PGF}\cap\mathcal{GF}_{n}\cap\mathcal{PGF}^{\perp}=(\mathcal{PGF}\cap\mathcal{PGF}^{\perp})\cap\mathcal{GF}_{n}=\mathcal{P}\cap\mathcal{GF}_{n}=\mathcal{P}.$
@@ 7.2. Application 5: Exact model structures on $\mathcal{GF}^{<\infty}$
Acknowledgement: We thank Ioannis Emmanouil for his useful comments.